# Late-Time Evolution of Magnetized Disks in Tidal Disruption Events

### v1

@@ 4.1 ASASSN–14li
ASASSN-14li is a TDE that discovered by the All-Sky Automated Survey for SuperNovae (ASAS-SN; Shappee et al. 2014; Holoien et al. 2016). It was first detected on 2014 November 22, in a post-starburst galaxy at a redshift of $z=0.0206$ (Jose et al., 2014). The late-time optical and UV light curves are well observed (Brown et al., 2017), showing gradual evolution that is consistent with emission from an accretion disk (van Velzen et al., 2019).
Since ASASSN-14li has been observed over much of the past decade, it serves as an ideal candidate for late-time light curve modeling. In this subsection, we analyze its late-time light curves [^6] Data taken from Mummery et al. (2024). (starting 380 days after its disruption) obtained with Swift/UVOT across five filters: B, U, UVM2, UVW1, and UVW2. Using the magnetized $\alpha$-viscosity model, we fit the data to estimate the Shakura-Sunyaev $\alpha$ parameter. For this analysis, we adopt a SMBH mass of $\log_{10}(M_{\bullet}/M_{\odot})=6.23^{+0.39}_{-0.40}$, based on velocity dispersion [^7] This mass estimate from galaxy scaling relations is statistically consistent with X-ray continuum fitting mass estimates (Wen et al., 2020), though the error bars on both are significant. measurements (Wevers et al., 2017).
We fit the observed light curves using those calculated from Eq. 4 with the self-similar solution. The free parameters in our model are the Shakura-Sunyaev $\alpha$ parameter and $\nu_{0,\rm S}$, which corresponds to the stellar mass. The total likelihood is computed as follows:
@@
ASASSN-14li’s viscous spreading is known to be slow (Wen et al., 2023); however, the fitted parameter $\alpha$ falls within a theoretically reasonable range (Hirose et al., 2006). There is therefore no need to invoke additional physics to further limit the spread of the disk outer edge, such as angular momentum loss in a magnetized wind (Wen et al., 2023).
@@ 4.2 Fossil TDEs
Even though the accretion disks of TDEs will eventually turn off, TDEs may remain bright in the UV for decades or centuries, allowing us to potentially detect fossil TDEs that erupted long ago. By fitting theoretical models to observations, we could measure parameters of the SMBHs, such as their mass, and thereby gain insights into their origins (Volonteri, 2010) and growth histories (Bhowmick et al., 2024), which often remain unclear. To accurately measure these parameters, however, we need to use the most precise models possible. Late-time TDE disks are much simpler phenomena than the early stages of a TDE, before an accretion disk has circularized, making it easier to develop theoretical models to fit observations. Therefore, it may be useful to build a large library of fossil TDEs.
The Ultraviolet Transient Astronomy Satellite (ULTRASAT), scheduled to be launched in 2027 (Shvartzvald et al., 2024), is a promising instrument for detecting these events. ULTRASAT is a near-UV space telescope ($230-290$nm), designed to conduct a wide-field UV survey for transients. The predicted TDE detection rate for ULTRASAT is enormous and will expand the current TDE sample by $1-2$ orders of magnitude (Shvartzvald et al., 2024), similar to the discovery potential of Rubin Observatory’s LSST but without the false positives that are likely to trouble a lower cadence optical survey (Bricman & Gomboc, 2020). Given this potential, we want to check whether ULTRASAT could also detect fossil TDEs. In this subsection, we focus on UV detection with ULTRASAT rather than optical surveys like LSST because fossils are brighter in the UV than in optical wavelengths. Furthermore, galaxies are much brighter in optical wavelengths, meaning that late-time fossils would likely be overwhelmed by background starlight, making them effectively undetectable in the optical.
@@
For detection, we assume a typical limiting AB magnitude of $m_{\rm AB}=22.5$ (Shvartzvald et al., 2024). Since TDEs can be observed in the UV up to cosmological distances, we account for K-corrections to determine the limiting luminosity distance for detecting fossil TDEs:
$D_{\rm L}^{\rm max}(z,t,M_{\bullet},M_{\star})=\sqrt{\frac{(1+z)L_{(1+z)\nu}(t,M_{\bullet},M_{\star})}{4\pi F_{\nu}}}$ (17)
where $L_{\nu}$ is the luminosity calculated from our model at the center of the ULTRASAT band ($250$nm), $F_{\nu}=10^{-\frac{m_{\rm AB+48.6}}{2.5}}\approx 3.6\times 10^{-29}\rm{\,erg\,s}^{-1}\,\rm{cm}^{-2}\,\rm{Hz}^{-1}$ is the limiting flux density of ULTRASAT, and $z$ is the redshift of the observed TDE. The limiting luminosity distance for a specific TDE depends on the redshift $z$, the TDE’s age $t$, the SMBH mass $M_{\bullet}$, and the stellar mass $M_{\star}$ involved. Additionally, for volume calculations we use the comoving distance:
@@
where $H_{0}=69.6\,\rm{km\,s^{-1}\,Mpc^{-1}}$ is the Hubble constant, and $\Omega_{\rm m}=0.286$ and $\Omega_{\Lambda}=0.714$ are the matter density and dark energy density, respectively (Bennett et al., 2014).
@@
Another constraint on the detectability of evolution is the photometric stability of ULTRASAT. This limit will likely allow us to detect luminosity changes of $\Delta L=1\%$ over a year (Shvartzvald et al., 2024), though we also consider a more conservative estimate [^8] This likely represents a worst-case scenario for ULTRASAT. Eran Ofek, private communication. of $\Delta L=5\%$.
@@
[…]
$\dot{n}_{\rm TDE}(M_{\bullet})\frac{dV(z,t)}{dz}\frac{dN_{\bullet}(z,M_{\bullet})}{dM_{\bullet}}\frac{dN_{\star}(M_{\star})}{dM_{\star}}$
×$\times\frac{\Delta\Omega}{4\pi}H(z,t,M_{\bullet},M_{\star})$
where we define the minimum age for a TDE to be classified as a fossil as $10$ years. $\dot{n}_{\rm TDE}$ is the volumetric TDE rate, which can be estimated from Zwicky Transient Facility (ZTF) observations (Yao et al., 2023) or theoretical estimations (Stone & Metzger, 2016). $V=\frac{4\pi}{3}D_{\rm C}^{3}$ is the comoving volume, $\frac{dN_{\bullet}}{dM_{\bullet}}$ is the SMBH mass function (Shankar et al., 2009), and $\frac{dN_{\star}}{dM_{\star}}$ is the Kroupa present-day mass function for an old nuclear population:
@@
To account for all the limiting factors described in the previous paragraph, we define a step function $H(z,t,M_{\bullet},M_{\star})$ that returns $1$ when the limiting conditions are satisfied and 00 otherwise.
The maximum SMBH mass, $M_{\rm Hills}(M_{\star})$, corresponds to the Hills mass (Hills, 1975), determined by the requirement that the tidal radius, $r_{t}$, must be outside the innermost bound circular orbit (IBCO). For simplicity, we use the Schwarzschild IBCO in our calculations. The maximum redshift, $z_{\rm max}(t,M_{\bullet},M_{\star})$, at which a fossil TDE can be detected is calculated using Eqs. 17 and 18. Furthermore, ULTRASAT will not survey the entire sky but only a fraction of it. In its low-cadence mode, ULTRASAT’s sky coverage is approximately $\Delta\Omega\approx 6800\,\rm{deg}^{2}$ (Shvartzvald et al., 2024).
Figure 7: Cumulative distribution functions showing the number of ULTRASAT-observable fossil TDEs as a function of their age (top) and their comoving distance and redshift (bottom). The limiting NUV AB magnitude in these estimations is $m_{\rm AB}=22.5$. The results are shown for a $3\sigma$ detection threshold for the minimum detectable luminosity change over 1 year, and with two different photometric stability limits: $\Delta L=1\%$ (blue and orange), and $\Delta L=5\%$ (green and pink). The predicted number of ULTRASAT fossils are presented for two different estimates of disk effective viscosity: $\alpha=0.01$ (orange and pink) and $\alpha=0.1$ (blue and green). In these calculations, we considered both the empirical TDE rate from the ZTF sample of TDEs (lower curves; Yao et al. 2023) and the dynamically predicted rate from loss cone theory (higher curves; Stone & Metzger 2016). Under our most pessimistic assumptions (low event rate, low $\alpha$, poor photometric stability), ULTRASAT will find $\sim$ few TDE fossils; under our most optimistic assumptions (high event rate, high $\alpha$, good photometric stability), $\sim 340$.
The results of our estimate for the number of observed fossil TDEs with $3\sigma$ detection are presented in Fig. 7, which shows the cumulative number of fossil TDEs as a function of their age and their distance. The results are presented for different ULTRASAT photometric stability thresholds, Shakura-Sunyaev $\alpha$ prefactor values, and TDE rates.
Additionally, by assuming less stringent detection limits, we can imagine building a low-specificity library of fossil candidates that could be verified through future observations with other instruments. Therefore, in Table 1, we present the total number of fossil TDEs and candidates that ULTRASAT could detect under different detection thresholds for luminosity evolution.
@@
$3\sigma$ 66 281 29 130
$2\sigma$ 120 515 48 215
$1\sigma$ 212 911 51 227
@@
$3\sigma$ 76 338 4 22
$2\sigma$ 151 669 6 35
$1\sigma$ 345 1484 7 38
Table 1: Total number of fossil TDEs and fossil candidates observable by ULTRASAT. We use the same parameters as in Fig. 7, but with different detection thresholds for the luminosity evolution ($3\sigma$, $2\sigma$, $1\sigma$). “Observations” and “Theory” refer to whether we use (low) observationally-motivated TDE rates or (high) theoretically-motivated TDE rates. In Table (a) we consider rapidly evolving disks ($\alpha=0.1$) and in Table (b) we consider slowly evolving ones ($\alpha=0.01$).
Fig. 7 and Table 1 show that for the $3\sigma$ detection threshold with the conservative threshold of $\Delta L=5\%$, the most pessimistic combination of assumptions ($\alpha=0.01$ and a TDE rate from ZTF) predicts the detection of a few fossil TDEs at $\sim 10$ years old, while the most optimistic scenario (with $\alpha=0.1$ and a TDE rate from theory) predicts the detection of $\sim 130$ fossil TDEs. For the less conservative threshold of $\Delta L=1\%$, ULTRASAT is expected to detect $\sim 70$ fossil TDEs from decades ago in the most pessimistic case (with $\alpha=0.1$ and a TDE rate from ZTF), and up to $\sim 340$ in the most optimistic case (with $\alpha=0.01$ and a TDE rate from theory).
There are qualitative differences between the low photometric stability ($\Delta L=5\%$) case and the high photometric stability ($\Delta L=1\%$) case. When $\Delta L$ is large, more fossils are detected as $\alpha$ increases, because the limiting factor for detection is observable time evolution. When $\Delta L$ is small, more fossils are detected as $\alpha$ decreases, because here the limiting factor for detection is UV flux. In the $\Delta L=5\%$ case, statistical uncertainties are generally unimportant, but they are dominant for $\Delta L=1\%$.
@@ 4.3 TDE-QPEs
Almost all galaxies above a certain size host a SMBH at their center, surrounded by a dense cluster of stars and compact objects. This dense nucleus is an excellent environment for various dynamical processes, such as TDEs, as well as the formation of extreme-mass ratio inspirals (EMRIs). An EMRI occurs when a stellar-mass compact object spirals toward the central SMBH through the emission of gravitational waves (GWs) (Amaro-Seoane, 2018). The compact object can approach very close to the SMBH while orbital energy is radiated away causing its semi-major axis to shrink. Recent studies suggest that the combination of a TDE with an EMRI may be the origin of a newly observed class of X-ray transients known as “quasi-periodic eruptions” (QPEs) (Linial & Metzger, 2023; Franchini et al., 2023).
QPEs are strong and short X-ray bursts that recur every few hours and originate near central SMBHs (Miniutti et al., 2019; Giustini et al., 2020; Arcodia et al., 2021; Chakraborty et al., 2021; Arcodia et al., 2024; Guolo et al., 2024; Nicholl et al., 2024). It is currently unknown what triggers these events and how long they last, but any viable explanation for the QPE phenomenon requires a mechanism capable of abruptly and quasi-periodically feeding the innermost region of a relatively low-mass SMBH.
One class of mechanisms for the QPEs are the collisions between an orbiting secondary object and an accretion disk around a primary SMBH (Suková et al., 2021; Xian et al., 2021; Linial & Metzger, 2023; Tagawa & Haiman, 2023; Franchini et al., 2023; Linial & Metzger, 2024). When an EMRI passes twice per orbit through the accretion disk it perturbs the disk and the disturbances can potentially trigger the observed QPEs. The QPE luminosity can be produced by shocked disk material ejection, bow shocks in the disk, or by enhanced accretion rate. Some of the most promising models suggest that the source of the accretion disk is from a TDE Linial & Metzger (2023); Franchini et al. (2023) and the interactions begin when the spreading disk from the TDE has sufficiently radially extended until it meets an EMRI.
This particular idea is supported by recent observations of a known TDE [^9] But see also Chakraborty et al. (2021); Guolo et al. (2025); Wevers et al. (2025). which has started to exhibit X-ray QPEs (Nicholl et al., 2024). The TDE AT2019qiz was first discovered by ZTF, and over several months, its UV and optical luminosity declined until reaching its plateau phase. Approximately $1500$ days after its first optical detection, the Chandra X-ray Observatory found repeating sharp increases in the X-ray luminosity. The mean recurrence time between peaks is $48.4\pm 0.3$ hours, with typical peak durations of $8-10$ hours, consistent with previous QPE observations.
@@
For simplicity, we assume a BH with mass $m_{\bullet}$ in a circular orbit as the secondary object in our calculations. The EMRI rate is highly uncertain (Babak et al., 2017), so we explore a range of possible values: $\dot{N}_{\rm EMRI}=10^{-6}-10^{-8}\,\rm{yr}^{-1}\rm{gal}^{-1}$ (Gair et al., 2004). Since the BH’s semi-major axis contracts due to the emission of GW, the EMRI rate is approximately $\dot{N}_{\rm EMRI}\approx 1/t_{\rm GW}$, where $t_{\rm GW}$ is the GW inspiral time: the time it takes the BH to evolve from an initial semi-major axis to merger. For a circular orbit with an initial semi-major axis $a_{0}$ the GW timescale is given by:
$t_{\rm GW}=\frac{5c^{5}a_{0}^{4}}{256G^{3}M_{\bullet}m_{\bullet}(M_{\bullet}+m_{\bullet})}.$ (21)
So, by knowing the EMRI rate, we can calculate the radius $r_{\rm sit}\equiv a_{0}$ where an EMRI is likely to be found. This radius corresponds to the radius that a TDE needs to spread in order to start exhibiting QPEs.
In Fig. 8, we show the time $t_{\rm on}$ it takes for the TDE to spread to the radius $r_{\rm sit}$ and the QPEs to turn on, as a function of the SMBH mass and the QPEs mean recurrence time. The results are shown for various parameter ranges of the EMRI rate, secondary mass, the alpha parameter, and the mass of the disrupted star. In all cases, the QPEs turn on before the time when Chandra observed the first QPE in AT2019qiz. Nicholl et al. (2024) were unable to constrain exactly when the QPEs began in AT2019qiz. Therefore, our results suggest that the QPEs may have started before the first observed X-ray burst, or alternatively could originate from a different mechanism. However, it is plausible that not all disk-orbiter collisions produce detectable X-ray flares, and additional conditions beyond the physical extent of the disk are required. The “TDE+EMRI=QPE” model could, in principle, be tested with more continuous X-ray monitoring of TDEs over a range timescales as shown in Fig. 8.
Figure 8: The time it takes for the QPEs to turn on after the TDE burst as a function of the SMBH mass (Top) and recurrence time (Bottom). The secondary object is a BH in a circular orbit with mass $m_{\bullet}$. The default parameters are $m_{\bullet}=10M_{\odot}$, $\dot{N}_{\rm EMRI}=10^{-7}\,\text{yr}^{-1}\text{gal}^{-1}$, $\alpha=0.1$, $M_{\bullet}=10^{6}M_{\odot}$, and $M_{\star}=0.3M_{\odot}$. Chandra observed the beginning of the QPEs $\sim 1500$ days after AT2019qiz was first detected, around a SMBH with $\log_{10}M_{\bullet}/M_{\odot}=6.3^{+0.3}_{-0.2}$, with a mean recurrence time between peaks of $48.4\pm 0.3$ hours (Nicholl et al., 2024). The observed time to exhibit QPEs of AT2019qiz is longer than the expected time from our model.

### v2

@@ 4.1 ASASSN–14li
ASASSN-14li is a TDE that discovered by the All-Sky Automated Survey for SuperNovae (ASAS-SN; Shappee+2014; Holoien+2016). It was first detected on 2014 November 22, in a post-starburst galaxy at a redshift of $z=0.0206$ (Jose+2014). The late-time optical and UV light curves are well observed (Brown+2017), showing gradual evolution that is consistent with emission from an accretion disk (vanVelzen+2019).
Since ASASSN-14li has been observed over much of the past decade, it serves as an ideal candidate for late-time light curve modeling. In this subsection, we analyze its late-time light curves [^6] Data taken from Mummery+2024. (starting 380 days after its disruption) obtained with Swift/UVOT across five filters: B, U, UVM2, UVW1, and UVW2. Using the magnetized $\alpha$-viscosity model, we fit the data to estimate the Shakura-Sunyaev $\alpha$ parameter. For this analysis, we adopt a SMBH mass of $\log_{10}(M_{\bullet}/M_{\odot})=6.23^{+0.39}_{-0.40}$, based on velocity dispersion [^7] This mass estimate from galaxy scaling relations is statistically consistent with X-ray continuum fitting mass estimates (Wen+2020), though the error bars on both are significant. measurements (Wevers+2017).
We fit the observed light curves using those calculated from Eq.˜4 with the self-similar solution. The free parameters in our model are the Shakura-Sunyaev $\alpha$ parameter and $\nu_{0,\rm S}$, which corresponds to the stellar mass. The total likelihood is computed as follows:
@@
ASASSN-14li’s viscous spreading is known to be slow (Wen+2023); however, the fitted parameter $\alpha$ falls within a theoretically reasonable range (Hirose+2006). There is therefore no need to invoke additional physics to further limit the spread of the disk outer edge, such as angular momentum loss in a magnetized wind (Wen+2023).
@@ 4.2 Fossil TDEs
Even though the accretion disks of TDEs will eventually turn off, TDEs may remain bright in the UV for decades or centuries, allowing us to potentially detect fossil TDEs that erupted long ago. By fitting theoretical models to observations, we could measure parameters of the SMBHs, such as their mass, and thereby gain insights into their origins (Volonteri2010) and growth histories (Bhowmick+2024), which often remain unclear. To accurately measure these parameters, however, we need to use the most precise models possible. Late-time TDE disks are much simpler phenomena than the early stages of a TDE, before an accretion disk has circularized, making it easier to develop theoretical models to fit observations. Therefore, it may be useful to build a large library of fossil TDEs.
The Ultraviolet Transient Astronomy Satellite (ULTRASAT), scheduled to be launched in 2027 (Shvartzvald+2024), is a promising instrument for detecting these events. ULTRASAT is a near-UV space telescope (230290230-290nm), designed to conduct a wide-field UV survey for transients. The predicted TDE detection rate for ULTRASAT is enormous and will expand the current TDE sample by 121-2 orders of magnitude (Shvartzvald+2024), similar to the discovery potential of Rubin Observatory’s LSST but without the false positives that are likely to trouble a lower cadence optical survey (Bricman&Gomboc2020). Given this potential, we want to check whether ULTRASAT could also detect fossil TDEs. In this subsection, we focus on UV detection with ULTRASAT rather than optical surveys like LSST because fossils are brighter in the UV than in optical wavelengths. Furthermore, galaxies are much brighter in optical wavelengths, meaning that late-time fossils would likely be overwhelmed by background starlight, making them effectively undetectable in the optical.
@@
For detection, we assume a typical limiting AB magnitude of $m_{\rm AB}=22.5$ (Shvartzvald+2024). Since TDEs can be observed in the UV up to cosmological distances, we account for K-corrections to determine the limiting luminosity distance for detecting fossil TDEs:
$D_{\rm L}^{\rm max}(z,t,M_{\bullet},M_{\star})=\sqrt{\frac{(1+z)L_{(1+z){\color[rgb]{1,0,0}\textbf{}}f}(t,M_{\bullet},M_{\star})}{4\pi F{\color[rgb]{1,0,0}\textbf{}}_{f}}}$ (17)
where $L_{f}$ is the luminosity calculated from our model at the center of the ULTRASAT band (250250nm), ${\color[rgb]{1,0,0}\textbf{}}F_{f}=10^{-\frac{m_{\rm AB+48.6}}{2.5}}\approx 3.6\times 10^{-29}\rm{\,erg\,s}^{-1}\,\rm{cm}^{-2}\,\rm{Hz}^{-1}$ is the limiting flux density of ULTRASAT, and $z$ is the redshift of the observed TDE. The limiting luminosity distance for a specific TDE depends on the redshift $z$, the TDE’s age $t$, the SMBH mass $M_{\bullet}$, and the stellar mass $M_{\star}$ involved. Additionally, for volume calculations we use the comoving distance:
@@
where $H_{0}=69.6\,\rm{km\,s^{-1}\,Mpc^{-1}}$ is the Hubble constant, and $\Omega_{\rm m}=0.286$ and $\Omega_{\Lambda}=0.714$ are the matter density and dark energy density, respectively (Bennett+2014).
@@
Another constraint on the detectability of evolution is the photometric stability of ULTRASAT. This limit will likely allow us to detect luminosity changes of $\Delta L=1\%$ over a year (Shvartzvald+2024), though we also consider a more conservative estimate [^8] This likely represents a worst-case scenario for ULTRASAT. Eran Ofek, private communication. of $\Delta L=5\%$.
@@
$N_{\rm fossils}$ $\approx\int_{0.08M_{\odot}}^{1M_{\odot}}dM_{\star}\int_{10^{5}M_{\odot}}^{M_{\rm Hills}}dM_{\bullet}\int_{10\,\rm years}^{200\,\rm years}dt\int_{0}^{z_{\rm max}}dz$ (19)
${\color[rgb]{1,0,0}\textbf{}}\dot{N}_{\rm TDE}(M_{\bullet})\frac{dV(z,t)}{dz}\frac{d{\color[rgb]{1,0,0}\textbf{}}n_{\bullet}(z,M_{\bullet})}{dM_{\bullet}}\frac{dN_{\star}(M_{\star})}{dM_{\star}}$
$\times\frac{\Delta\Omega}{4\pi}H(z,t,M_{\bullet},M_{\star})$
where we define the minimum age for a TDE to be classified as a fossil as $10$ years. Even though TDEs remain bright for many centuries, identifying them as fossil TDEs is more challenging because their classification also requires observable time evolution. Therefore, we limit our integration to $200$ years, which, as shown in our results in Fig.˜7, is a reasonable choice. $\dot{N}_{\rm TDE}$ is the per-galaxy TDE rate, which can be estimated from Zwicky Transient Facility (ZTF) observations ($\dot{N}_{\rm TDE}(M_{\bullet})=4.1\times 10^{-5}\left(\frac{M_{\bullet}}{10^{6}M_{\odot}}\right)^{-0.25}$ yr-1 gal-1; Yao+2023) or theoretical estimations ($\dot{N}_{\rm TDE}(M_{\bullet})=1.9\times 10^{-4}\left(\frac{M_{\bullet}}{10^{6}M_{\odot}}\right)^{-0.404}$ yr-1 gal-1; Stone+2016). $V=\frac{4\pi}{3}D_{\rm C}^{3}$ is the comoving volume, $\frac{d{\color[rgb]{1,0,0}\textbf{}}n_{\bullet}}{dM_{\bullet}}$ is the volumetric SMBH mass function (Shankar+2009), and $\frac{dN_{\star}}{dM_{\star}}$ is the Kroupa present-day mass function for an old nuclear population:
@@
To account for all the limiting factors described in the previous paragraph, we define a step function $H(z,t,M_{\bullet},M_{\star})$ that returns $1$ when the limiting conditions are satisfied and 0 otherwise.
The maximum SMBH mass, $M_{\rm Hills}(M_{\star})$, corresponds to the Hills mass (Hills1975), determined by the requirement that the tidal radius, $r_{t}$, must be outside the innermost bound circular orbit (IBCO). For simplicity, we use the Schwarzschild IBCO in our calculations. We use $10^{5}M_{\odot}$ as the lower integration limit for the SMBH mass, as this is the smallest value in the SMBH mass function from Shankar+2009. Additionally, detectable fossil TDEs are limited by their luminosity, which increases with SMBH mass, as shown in Fig.˜6.
The maximum redshift, $z_{\rm max}(t,M_{\bullet},M_{\star})$, at which a fossil TDE can be detected is calculated using Eqs.˜17 and 18. Furthermore, ULTRASAT will not survey the entire sky but only a fraction of it. In its low-cadence mode, ULTRASAT’s sky coverage is approximately $\Delta\Omega\approx 6800\,\rm{deg}^{2}$ (Shvartzvald+2024).
Figure 7: Cumulative distribution functions showing the number of ULTRASAT-observable fossil TDEs as a function of their age (top) , their comoving distance and redshift (middle) and SMBH mass (bottom). The limiting NUV AB magnitude in these estimations is $m_{\rm AB}=22.5$. The results are shown for a $3\sigma$ detection threshold for the minimum detectable luminosity change over 1 year, and with two different photometric stability limits: $\Delta L=1\%$ (blue and orange), and $\Delta L=5\%$ (green and pink). The predicted number of ULTRASAT fossils are presented for two different estimates of disk effective viscosity: $\alpha=0.01$ (orange and pink) and $\alpha=0.1$ (blue and green). In these calculations, we considered both the empirical TDE rate from the ZTF sample of TDEs (lower curves; Yao+2023) and the dynamically predicted rate from loss cone theory (higher curves; Stone+2016). Under our most pessimistic assumptions (low event rate, low $\alpha$, poor photometric stability), ULTRASAT will find $\sim$ few TDE fossils; under our most optimistic assumptions (high event rate, high $\alpha$, good photometric stability), $\sim 210$.
The results of our estimate for the number of observed fossil TDEs with $3\sigma$ detection are presented in Fig.˜7, which shows the cumulative number of fossil TDEs as a function of their age and their distance. The results are presented for different ULTRASAT photometric stability thresholds, Shakura-Sunyaev $\alpha$ prefactor values, and TDE rates.
Additionally, by assuming less stringent detection limits, we can imagine building a low-specificity library of fossil candidates that could be verified through future observations with other instruments. Therefore, in Table˜1, we present the total number of fossil TDEs and candidates that ULTRASAT could detect under different detection thresholds for luminosity evolution.
@@
$3\sigma$ 34 147 17 74
$2\sigma$ 58 250 25 108
$1\sigma$ 93 396 26 111
@@
$3\sigma$ 47 207 4 25
$2\sigma$ 92 401 7 40
$1\sigma$ 196 833 8 43
Table 1: Total number of fossil TDEs and fossil candidates observable by ULTRASAT. We use the same parameters as in Fig.˜7, but with different detection thresholds for the luminosity evolution ($3\sigma$, $2\sigma$, $1\sigma$). “Observations” and “Theory” refer to whether we use (low) observationally-motivated TDE rates or (high) theoretically-motivated TDE rates. In Table (a) we consider rapidly evolving disks ($\alpha=0.1$) and in Table (b) we consider slowly evolving ones ($\alpha=0.01$).
Fig.˜7 and Table˜1 show that for the $3\sigma$ detection threshold with the conservative threshold of $\Delta L=5\%$, the most pessimistic combination of assumptions ($\alpha=0.01$ and a TDE rate from ZTF) predicts the detection of a few fossil TDEs at $\sim 10$ years old, while the most optimistic scenario (with $\alpha=0.1$ and a TDE rate from theory) predicts the detection of 75 fossil TDEs. For the less conservative threshold of $\Delta L=1\%$, ULTRASAT is expected to detect 35 fossil TDEs from decades ago in the most pessimistic case (with $\alpha=0.1$ and a TDE rate from ZTF), and up to 210 in the most optimistic case (with $\alpha=0.01$ and a TDE rate from theory).
There are qualitative differences between the low photometric stability ($\Delta L=5\%$) case and the high photometric stability ($\Delta L=1\%$) case. When $\Delta L$ is large, more fossils are detected as $\alpha$ increases, because the limiting factor for detection is observable time evolution. When $\Delta L$ is small, more fossils are detected as $\alpha$ decreases, because here the limiting factor for detection is UV flux. In the $\Delta L=5\%$ case, statistical uncertainties are generally unimportant, but they are dominant for $\Delta L=1\%$. Additionally, in all cases, we find that most detectable fossil TDEs originate from low-mass SMBHs, indicating that the limiting factor for detectability is the evolution rate rather than the source luminosity.
@@ 4.3 TDE-QPEs
Almost all galaxies above a certain size host a SMBH at their center, surrounded by a dense cluster of stars and compact objects. This dense nucleus is an excellent environment for various dynamical processes, such as TDEs, as well as the formation of extreme-mass ratio inspirals (EMRIs). An EMRI occurs when a stellar-mass compact object spirals toward the central SMBH through the emission of gravitational waves (GWs) (Amaro-Seoane2018). The compact object can approach very close to the SMBH while orbital energy is radiated away causing its semi-major axis to shrink. Recent studies suggest that the combination of a TDE with an EMRI may be the origin of a newly observed class of X-ray transients known as “quasi-periodic eruptions” (QPEs) (Linial&Metzger2023; Franchini+2023).
QPEs are strong and short X-ray bursts that recur every few hours and originate near central SMBHs (Miniutti+2019; Giustini+2020; Arcodia+2021; Chakraborty+2021; Arcodia+2024; Guolo+2024; Nicholl+2024). It is currently unknown what triggers these events and how long they last, but any viable explanation for the QPE phenomenon requires a mechanism capable of abruptly and quasi-periodically feeding the innermost region of a relatively low-mass SMBH.
One class of mechanisms for the QPEs are the collisions between an orbiting secondary object and an accretion disk around a primary SMBH (Sukova+2021; Xian+2021; Linial&Metzger2023; Tagawa&Haiman2023; Franchini+2023; Linial&Metzger2024). When an EMRI passes twice per orbit through the accretion disk it perturbs the disk and the disturbances can potentially trigger the observed QPEs. The QPE luminosity can be produced by shocked disk material ejection, bow shocks in the disk, or by enhanced accretion rate. Some of the most promising models suggest that the source of the accretion disk is from a TDE Linial&Metzger2023; Franchini+2023 and the interactions begin when the spreading disk from the TDE has sufficiently radially extended until it meets an EMRI.
This particular idea is supported by recent observations of a known TDE [^9] But see also Chakraborty+2021; Guolo+2025; Wevers+2025. which has started to exhibit X-ray QPEs (Nicholl+2024). The TDE AT2019qiz was first discovered by ZTF, and over several months, its UV and optical luminosity declined until reaching its plateau phase. Approximately $1500$ days after its first optical detection, the Chandra X-ray Observatory found repeating sharp increases in the X-ray luminosity. The mean recurrence time between peaks is $48.4\pm 0.3$ hours, with typical peak durations of 8108-10 hours, consistent with previous QPE observations.
@@
For simplicity, we assume a BH with mass $m_{\bullet}$ in a Keplerian circular orbit as the secondary object in our calculations. If the secondary object crosses the accretion disk twice per orbit, the QPE recurrence time is given by $t_{\rm rec}=\pi\sqrt{\frac{r_{\rm sit}^{3}}{GM_{\bullet}}}$ where $r_{\rm sit}$ is the radius at which the secondary object is likely to be located. This radius corresponds to the radius that a TDE needs to spread in order for QPEs to begin. By measuring $t_{\rm rec}$ we can calculate $r_{\rm sit}$ and then calculate $t_{\rm on}$, the time it takes for the TDE to spread to the radius $r_{\rm sit}$ and the QPEs to turn on.
In Fig.˜8, we show the time $t_{\rm on}$ as a function of the SMBH mass and the QPEs mean recurrence time. The results are shown for various parameter ranges of the secondary mass, the alpha parameter, and the mass of the disrupted star. In all cases, the QPEs turn on before the time when Chandra observed the first QPE in AT2019qiz. Nicholl+2024 were unable to constrain exactly when the QPEs began in AT2019qiz. Therefore, our results suggest that the QPEs may have started before the first observed X-ray burst, or alternatively could originate from a different mechanism. However, it is plausible that not all disk-orbiter collisions produce detectable X-ray flares, and additional conditions beyond the physical extent of the disk are required. The “TDE+EMRI=QPE” model could, in principle, be tested with more continuous X-ray monitoring of TDEs over a range timescales as shown in Fig. 8.
Figure 8: The time it takes for the QPEs to turn on after the TDE burst as a function of the SMBH mass (Top) and recurrence time (Bottom). The secondary object is a BH in a circular orbit with mass $m_{\bullet}$. The default parameters are $m_{\bullet}=10M_{\odot}$, $t_{\rm rec}=48.4$ hours, $\alpha=0.1$, $M_{\bullet}=10^{6}M_{\odot}$, and $M_{\star}=0.3M_{\odot}$. Chandra observed the beginning of the QPEs $\sim 1500$ days after AT2019qiz was first detected, around a SMBH with $\log_{10}M_{\bullet}/M_{\odot}=6.3^{+0.3}_{-0.2}$, with a mean recurrence time between peaks of $48.4\pm 0.3$ hours (Nicholl+2024). The observed time to exhibit QPEs of AT2019qiz is longer than the expected time from our model.